OpenStudy (anonymous):

3√(2x)=12. solve for x

OpenStudy (luigi0210):

Get the square root alone. So first divide that 3 out. Then square and divide again.

OpenStudy (anonymous):

what do you mean by that?

OpenStudy (kirbykirby):

The goal is just to get "x" by itself. \(3\sqrt{2x}=12\) \((3\sqrt{12x})^2=12^2\) this will get rid of the square root \(9(2x)=144\) \(18x=144\) \(x=\frac{144}{18}=8\)

OpenStudy (kirbykirby):

oops on the second step I made a typo: \((3\sqrt{2x})^2=12^2\)

OpenStudy (anonymous):

thank you so much

OpenStudy (luigi0210):

That if you want it done for ya ;)

OpenStudy (anonymous):

wait. @kirbykirby how did u get 9(2x)

OpenStudy (kirbykirby):

Recall the rule \((a\cdot b)^2=a^2\cdot b^2\) Here, you have \((3\sqrt{2x})^2\), so think of \(a=3\) and \(b=\sqrt{2x}\) Thus, you get \((3\sqrt{2x})^2=3^2\cdot (\sqrt{2x})^2\)

OpenStudy (anonymous):

yes but wouldn't the (√2x)^2 be √2x^2 ?

OpenStudy (kirbykirby):

\(\sqrt{2x}\ne \sqrt{2}x\) do you see the difference? You wrote √(2x) which I assumed you meant \(\sqrt{2x}\)

OpenStudy (kirbykirby):

If you have trouble convincing yourself. Recall this: \(\sqrt{x}\sqrt{y}=\sqrt{x y}\) Thus, \((\sqrt{2x})^2=(\sqrt{2}\sqrt{x})^2=(\sqrt{2})^2(\sqrt{x})^2=2x\)

OpenStudy (kirbykirby):

the third step is an application of the rule \((a\cdot b)^2=a^2\cdot b^2\) where \(a=\sqrt{2}\) and \(b=\sqrt{x}\)

OpenStudy (kirbykirby):

but really, it's just easier to think of anything squared, under a square root, removed the square root. So even if you have a complicated expression under the square root, you can just remove it. For example: \[\left(\sqrt{x+y^{1/2}+zx+\frac{2}{\pi}}\right)^2=x+y^{1/2}+zx+\frac{2}{\pi}\]

OpenStudy (kirbykirby):

Hopefully I didn't scare you with that last expression ^ :)

OpenStudy (anonymous):

lol its okay. thank u

OpenStudy (anonymous):

i just figured it out. u simplified it

OpenStudy (kirbykirby):

:)